Evaluate all possible principal angles in standard position
sinE=1

Answer by eperette(173)

Concepts:
sinE=y/r
cosE=x/r
tanE=y/x
cscE=r/y
secE=r/x
cotE=x/y
x^2 + y^2 = r^2
Since sinE=1=1/1
Then y=1 and r=1
x^2 + y^2 = r^2
x^2 + (1)^2 = (1)^2
x^2 + 1 = 1
x^2 = 0
Thus
sinE=1
cosE=x/r=0/1=0
tanE=y/x=1/0= undefined
cscE=1
secE=1/0 = undefined
cotE=0/1 = 0
*******THIS ONE IS ASKING FOR ALL POSSIBLE PRINCIPAL ANGLES...JUST REALIZE I DID PRINCIPAL RATIOS****
E= sin^-1(1)= 90 degrees